Let U = {1, 2, 3, 4, 5, 6, 7}, A = {2, 4, 6}, B = {3, 5} and C = {1, 2, 4, 7}, find

(i) A′ ∪ (B ∩ C′)

(ii) (B – A) ∪ (A – C)

Grade:12

1 Answers

Harshit Singh
3 years ago
Welcome to AskIITians

Given,

U = {1, 2, 3, 4, 5, 6, 7}, A = {2, 4, 6}, B = {3, 5} and C = {1, 2, 4, 7}

(i) A′ = {1, 3, 5, 7}

C′ = {3, 5, 6}

B ∩ C′ = {3, 5}

A′ ∪ (B ∩ C′) = {1, 3, 5, 7}

(ii) B – A = {3, 5}

A – C = {6}

(B – A) ∪ (A – C) = {3, 5, 6}

Thanks